OpenStudy (anonymous):

P(-6,5), 7y+4x=2 what it the perpendicular line for this?

OpenStudy (anonymous):

first y=mx+b slope intercept form y=-4/7x+2/7 then point slope form, change -4/7 to 7/4 y-5=7/4x+6 y=7/4+11